OpenStudy (anonymous):

Can someone tell me how to solve 4*(x+2)=14?? Step by Step? Slowly?

OpenStudy (anonymous):

4*(x+2) =14 That means 4(x+2) =14 The bracket means multiply so you multiply 4 x x =4x then 4 x 2 =8 so you have 4x + 8 = 14 so your trying to find the value of 'x' so you put 4x = 14 when a positive figure goes over the other side ' = sign it becomes a negative so its 4x = 14-8 so 4x = 6 so x = 6/4 a multiplication becomes a division when you take it over the equal sign which is '4' so x= 6/4 therefore 6/4 is cancelled in its lowest term so x =3/2 or x = 1 1/2